OpenStudy (anonymous):

A Realtor earns a commission of 5.5% on the sale of a house. If the commission the Realtor earned was $22,000, for how much was the house sold? Will I have to make an equation? And if so, what should it be?

OpenStudy (anonymous):

Yes, it will be 5.5/100=22,000/x 5.5x=2200000 x=400,000 The house was sold for 400,000

OpenStudy (anonymous):

@Veelution so its 5.5 divided by 100 = 22,000 divided by x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@veelution what do I do to get 5.5x = 200,000? What happens to the 100?

OpenStudy (phi):

5.5% of Sales Price = 22,000 let x be the sales price \[ \frac{5.5}{100} x = 22000 \] multiply both sides by 100 \[ \cancel{100}\cdot \frac{5.5}{\cancel{100}} x= 2,200,000 \]

OpenStudy (phi):

to find x, divide both sides by 5.5 if you do not like decimals, first multiply both sides by 10 to get 55 x = 22,000,000 (both sides got multiplied by 10, so 5.5 becomes 55) now divide by 55 on both sides x= 22000000/55 (22 million divided by 55)

OpenStudy (phi):

Here are some videos on percent that might be useful http://www.khanacademy.org/math/arithmetic/percents/v/describing-the-meaning-of-percent